### v1

# On quivers satisfying Quantum Yang-Baxter equation, and Hecke condition

Abstract.
In this paper we initiate the study of quivers satisfying quantum Yang-Baxter equation. By considering quivers with quantum group like relations such as quantum Yang-Baxter equation, Hecke condition, we follow RTT construction to produce examples of Leavitt path algebras that contain quantum matrix algebra as subalgebra. Furthermore, We show that the quantum Yang-Baxter equation, and Hecke condition for our RTT construction can be generated intrinsically from the adjacency matrix of certain quivers.
1. Introduction
The Quantum Yang-Baxter Equation (QYBE) is a fundamental consistency condition that arises in mathematical physics, quantum groups, and noncommutative algebra. Its origin traces back to influential papers of Yang [15] on quantum mechanical many-body problem and Baxter [3] on solution of the eight vertex model. A solution to the quantum Yang-Baxter equation is a linear map $R:V\otimes V\rightarrow V\otimes V$ where $V$ is a vector space over a field such that the equality $(R\otimes I_{V})(I_{V}\otimes R)(R\otimes I_{V})=(I_{V}\otimes R)(R\otimes I_{V})(I_{V}\otimes R)$ of operators holds in the space $\operatorname{End}(V\otimes V\otimes V)$. The matrix of such a linear map is called an $R$-matrix.
The term Quantum Yang–Baxter Equation (QYBE) was coined by Faddeev in the late 1970s. Intuitively, the QYBE ensures that when three interacting particles or algebraic objects are rearranged, the final outcome does not depend on the order in which the pairwise interactions occur. It has applications in the study of exactly solvable models in statistical mechanics and quantum field theory, and has connections to knot theory, braid theory and Hopf algebras. In [7] Drinfeld suggested to study the set-theoretical solutions of QYBE. A set-theoretical solution is a solution for which $V$ is a vector space spanned by a set $X$, and $R$ is the linear operator induced by a mapping $X\times X\rightarrow X\times X$. For geometric and algebraic interpretations of such solutions, the reader is referred to Etingof et al. [9].
In Section 2 of this paper we introduce the notion of quiver satisfying the quantum Yang-Baxter equation. Our approach is quite different from the one in [2]. We say that a quiver $Q$ satisfies the QYBE if the adjacency matrix of the Kronecker square quiver satisfies the QYBE. We show that if a quiver $Q$ satisfies the QYBE then its adjacency matrix $A$ satisfies the “Temperley-Lieb idempotent” property, that is, $A^{2}=\mu A$ for some scalar $\mu$. In particular, we show that a finite symmetric quiver $Q$ satisfies the QYBE if and only if each nontrivial connected component of $Q$ is a complete weighted quiver (that is, there is a nonzero weighted arrow between any two vertices in the component).
Closely linked to the QYBE is the Hecke condition, which places an additional polynomial constraint on the $R$-matrix. In its common form, the Hecke condition states that $(R-qI)(R+q^{-1}I)=0$ for some nonzero parameter $q$. Together, the QYBE and the Hecke condition provide a powerful algebraic framework used to construct quantum symmetries and invariants. Solutions that satisfy both conditions lead to braided monoidal categories, quantum knot invariants like the Jones polynomial, and integrable lattice models in statistical mechanics. In noncommutative algebra, these equations guide the formation of deformed algebras where commutation rules are governed by the $R$-matrix. In Section 3 of this paper, we give a recipe to construct Hecke $R$-matrices from the weighted adjacency matrices of certain quivers.
Recently, there has been a lot of interest in quantum graphs among researchers working in quantum communication theory. Quantum graphs were first introduced in 1930s by Pauling to describe organic molecules and since then they have been studied from different perspectives. Physicists define it as a metric graph with the Schrödinger operator defined at the edges (according to the physical motivation) and some coupling conditions at the graph vertices that specify the transition probabilities across the vertices (see for example [8]). Weaver [13] proposed quantum graphs as reflexive symmetric quantum relations on a von Neumann algebra. In [4], quantum graphs are defined in terms of a self-adjoint operator on $L^{2}(H)$, for a finite-dimensional $C^{*}$-algebra $H$, which is idempotent under pointwise multiplication generalizing the usual adjacency matrix of a graph. Brannan at al adapt this notion in [5] to directed graphs and develop the notion of quantum graph $C^{*}$-algebras. Leavitt path algebras are algebraic analogues of graph C*-algebras defined as certain quotients of path algebras where the relations are inspired from Cuntz-Krieger relations for graph $C^{*}$-algebras [1]. Leavitt path algebras have deep connections with symbolic dynamics and since they capture the homological properties of both path algebras and their Koszul dual, they form an important class of noncommutative algebras. It is known that the Leavitt path algebra over Dynkin quiver $A_{n}$ is the matrix algebra $M_{n}(\mathbb{K})$ and so it is natural to ask if quantum matrix algebra can be realized as Leavitt path algebra over some quiver.
In Section 4, we consider quivers with RTT relations determined by Hecke R-matrices and quantum Yang-Baxter equation and obtain quantum matrix algebra as subalgebra of the Leavitt path algebras over quivers with such relations. There is a natural operator-algebraic interpretation of the RTT-enhanced Leavitt path algebras constructed in this section. The imposition of local RTT relations at vertices with loop spaces replaces the free $C^{*}$-correspondence of [6] by a braided tensor correspondence determined by the Hecke $R$-matrix $R_{u}$. In the classical setting ($R_{u}=\mathrm{Flip}$), the correspondence is symmetric and yields the usual graph $C^{*}$-algebra. When $R_{u}$ is nontrivial, the tensor product over loop space $L(u)$ becomes a braided tensor product, placing $L_{\mathbb{K}}(Q,R)$ into the framework of Hayashi’s face-type quantum group correspondences [10] and more general quantum graph $C^{*}$-algebras. In particular, for a vertex $u$ with $\dim_{\mathbb{K}}L(u)=n^{2}$ and Hecke $R$-matrix arising from $GL_{q}(n)$, the subalgebra generated by the loop edges yields the quantum matrix algebra $\mathcal{O}_{q}(M_{n})$, and the resulting $L_{\mathbb{K}}(Q,R)$ may be viewed as a quantum graph algebra whose quantum symmetries are encoded combinatorially by the quiver and the RTT relations.
Thus, the construction developed here produces natural algebraic models for quantum Cuntz–Krieger algebras and braided graph correspondences, enriching the interface between Leavitt path algebras, quantum groups, and noncommutative topology.
2. Quivers Satisfying the QYBE
Definition 2.1.
Let $Q$ be a quiver (that is, a directed graph) with the adjacency matrix $A=[a_{ij}]$. The Kronecker square $\widehat{Q}$ is a quiver whose adjacency matrix is $A\otimes A$ where $A\otimes A$ is the Kronecker product matrix given as $A\otimes A=[b_{ij}]$ with $b_{ij}=a_{ij}A$.
[…]
$\widehat{Q_{0}}=\{[v_{i},v_{j}]\,:\,v_{i},v_{j}\in Q_{0}\}$
$\widehat{Q_{1}}=\{[e_{i},e_{j}]\,:\,e_{i},e_{j}\in Q_{1}\}$
$\qquad\widehat{s}([e,f])=[s(e),s(f)]$
$\qquad\widehat{r}([e,f])=[r(e),r(f)]$
Recall that an $n^{2}\times n^{2}$ matrix $X$ is said to satisfy the quantum Yang-Baxter Equation (QYBE) if
$(X\otimes I_{n})(I_{n}\otimes X)(X\otimes I_{n})=(I_{n}\otimes X)(X\otimes I_{n})(I_{n}\otimes X)$
where $I_{n}$ is the $n\times n$ identity matrix.
Definition 2.2.
Let $Q$ be a quiver with adjacency matrix $A$. We will say that $Q$ satisfies the quantum Yang-Baxter Equation if the adjacency matrix $X=A\otimes A$ of $\widehat{Q}$ satisfies the quantum Yang-Baxter Equation.
We first investigate when an adjacency matrix $X=A\otimes A$ satisfies the QYBE. It will quickly be apparent that such $X$ have a highly degenerate structure. This means $X$ will generally not be an $R$-matrix as it will be singular (barring the identity matrix); however, in the next section, we will use such $X$ to produce Hecke $R$-matrices, $\check{R}(q)$ which arise in a straightforward, combinatorial way via weighted adjacency matrices of complete quivers.
A quick computation introduces the equation of interest for this section. Suppose we have a matrix $Y=A\otimes B\in\mathbb{C}^{n^{2}\times n^{2}}$ where $A,B\in C^{n\times n}$. We want to determine when $Y$ satisfies the QYBE. We must have
$\qquad\qquad\qquad\qquad\,\,\,\,\,(Y\otimes I_{n})(I_{n}\otimes Y)(Y\otimes I_{n})=(I_{n}\otimes Y)(Y\otimes I_{n})(I_{n}\otimes Y)$
$\iff(A\otimes B\otimes I_{n})(I_{n}\otimes A\otimes B)(A\otimes B\otimes I_{n})=(I_{n}\otimes A\otimes B)(A\otimes B\otimes I_{n})(I_{n}\otimes A\otimes B)$
$\iff\qquad\qquad\quad\,(A\otimes BA\otimes B)(A\otimes B\otimes I_{n})=(A\otimes AB\otimes B)(I_{n}\otimes A\otimes B)$
$\iff\qquad\qquad\qquad\qquad\qquad\,\,\,A^{2}\otimes BAB\otimes B=A\otimes ABA\otimes B^{2}$
In the case where $X=A\otimes A$, the above equation becomes
(1) $A^{2}\otimes A^{3}\otimes A=A\otimes A^{3}\otimes A^{2}.$
The main theorem of this section, Theorem 2.3, characterizes the matrices $A$ that satisfy equation (1) above and finds that such matrices have the “Temperley-Lieb idempotent” property, that is, $A^{2}=\mu A$ for some scalar $\mu$.
Theorem 2.3.
Let $A$ be a $n\times n$ matrix, such that $A^{3}\neq 0$, then
$A\otimes A^{3}\otimes A^{2}=A^{2}\otimes A^{3}\otimes A$
if and only if $A^{2}\otimes A=A\otimes A^{2}$ if and only if $A^{2}=\mu A$.
Theorem 2.3 is an immediate consequence of Lemma 2.4, Lemma 2.5 and Lemma 2.6 below.
Lemma 2.4.
Suppose $X,Y,U,V$ are nonzero square matrices of the same size, then if $X\otimes Y=U\otimes V$, then there exists a nonzero scalar $\lambda$ such that $X=\lambda U$ and $Y=\frac{1}{\lambda}V$.
Proof.
Let $X,Y,U,V$ be $n\times n$ matrices over $\mathbb{R}$ (a similar argument works for $\mathbb{C}$). Define
$x=\text{vec}(X),\quad y=\text{vec}(Y),\quad u=\text{vec}(U),\quad v=\text{vec}(V).$
The equality $X\otimes Y=U\otimes V$ thus corresponds to the equality $x\otimes y=u\otimes v$. Note, this is not to say $\text{vec}(X\otimes Y)=x\otimes y$, but rather, if $x\otimes y=u\otimes v$, then the entries of $X\otimes Y$ must be the same as the entries of $U\otimes V$, so it suffices to show $x\otimes y=u\otimes v$. Furthermore, the equality $x\otimes y=u\otimes v$ is equivalent to the equality $xy^{T}=uv^{T}$ of rank one matrices.
As a rank one matrix, the singular value decomposition of $xy^{T}$ can be written as $xy^{T}=\sigma pq^{T}$, where
[…]
Similarly, one can write $uv^{T}=\tau rs^{T}$ where
[…]
As $xy^{T}=uv^{T}$, we must have $\sigma pq^{T}=\tau rs^{T}$.
For a nonzero rank one matrix, the left and right singular values are unique up to sign, thus there is a scalar $\alpha$ with
$p=\alpha r,\quad q=\frac{1}{\alpha}s,\quad\sigma=\tau.$
[…]
As $x=\lambda u$ and $y=\lambda^{-1}v$, we must have $X=\lambda U$ and $Y=\frac{1}{\lambda}V$, as desired. ∎
Lemma 2.5.
Let $A$ be an $n\times n$ matrix and suppose $A^{3}\neq 0$, then $A^{2}\otimes A=A\otimes A^{2}$ if and only if $A^{2}\otimes A^{3}\otimes A=A\otimes A^{3}\otimes A^{2}$.
Proof.
Suppose $A^{2}\otimes A=A\otimes A^{2}$. As $A^{2}\otimes A=A\otimes A^{2}$, we must have
$(A\otimes A)(A^{2}\otimes A)$ $=(A\otimes A)(A\otimes A^{2})$
$\Rightarrow A^{3}\otimes A^{2}$ $=A^{2}\otimes A^{3}$
Furthermore, if $I$ is the $n\times n$ identity matrix, then we find
$(A\otimes I)(A^{2}\otimes A)$ $=(A\otimes I)(A\otimes A^{2})$
$\Rightarrow A^{3}\otimes A$ $=A^{2}\otimes A^{2}$
and similarly, $A^{2}\otimes A^{2}=A\otimes A^{3}$. All together, we have the equality
$A\otimes A^{3}=A^{2}\otimes A^{2}=A^{3}\otimes A.$
Using the above equalities, we get the equation
$A^{2}\otimes A^{3}\otimes A=A^{3}\otimes A^{2}\otimes A=A^{3}\otimes A\otimes A^{2}=A\otimes A^{3}\otimes A^{2}.$
For the converse, we write
$x=\text{vec}(X),\qquad y=\text{vec}(A^{3})\qquad z=\text{vec}(A^{2}).$
The equality $A\otimes A^{3}\otimes A^{2}=A^{2}\otimes A^{3}\otimes A$ is equivalent to the equality $x\otimes y\otimes z=z\otimes y\otimes x$. As $A^{3}\neq 0$, we have $x,y,z\neq 0$. Further, the uniqueness of a pure tensor decomposition for three non-zero factors implies there exists scalars $\alpha,\beta,\gamma$ with
$x=\alpha z,\quad y=\beta y,\quad z=\gamma x,\quad\alpha\beta\gamma=1.$
As $\beta=1$, we get $\alpha\gamma=1$. As such, after translating the vectorized equalities back into square matrices, we find $A=\alpha A^{2}$ and $A^{2}=\gamma A$. It follows
$A^{2}\otimes A=(\gamma A)\otimes A=A\otimes(\gamma A)=A\otimes A^{2}.$
Lemma 2.6.
Let $A$ be a nonzero $n\times n$ matrix, then $A^{2}\otimes A=A\otimes A^{2}$ if and only if $A^{2}=\mu A$ for $\mu$ a scalar.
Proof.
The converse was shown in the above proof, so we prove sufficiency. If $A^{2}\otimes A=A\otimes A^{2}$, then by Lemma 2.4, assuming $A^{2}\neq 0$, there exists $\lambda\neq 0$ with $A=\lambda A^{2}$ and $A^{2}=\frac{1}{\lambda}A$. Letting $\mu=\frac{1}{\lambda}$, the result follows. ∎
Proof of Theorem 2.3. By combining Lemma 2.5 and Lemma 2.6, we have
$A^{2}\otimes A^{3}\otimes A=A\otimes A^{3}\otimes A^{2}\quad\iff\qquad A^{2}\otimes A=A\otimes A^{2}\quad\iff\qquad A^{2}=\mu A$
for $\mu$ a scalar.
We now restrict our focus to adjacency matrices of a graph $\Gamma$. Equivalently, we can interpret these graph adjacency matrices as adjacency matrices of a quiver $Q$ with bidirectional arrows.
In this section, we assume $A^{3}\neq 0$. From Theorem 2.3, it suffices to find graphs whose adjacency matrix satisfies $A^{2}=\mu A$ for some scalar $\mu$. We consider three different cases: The case where $\Gamma$ is without loops and multiple edges (Proposition 2.7), the case where $\Gamma$ has loops but no multiple edges (Proposition 2.8) and lastly, when $\Gamma$ has loops and multiple edges (Theorem 2.9).
Proposition 2.7.
Suppose $A$ is an adjacency $n\times n$ matrix corresponding to a graph $\Gamma$ without any loops or multiple edges, then $A^{2}=\mu A$ if and only if $A$ is the zero matrix, i.e. $\Gamma$ consists entirely of isolated vertices.
Proof.
As $\Gamma$ does not have any loops, we have $A_{ii}=0$ for all $i$, further each $A_{ij}$ is either 0 or $1$. By Theorem 2.3, it follows
$0=\mu A_{ii}=(A^{2})_{ii}=\sum_{j=1}^{n}A_{ij}A_{ji}=\sum_{j=1}^{n}A_{ij}=\text{deg}(i).$
We have thus shown $\deg(i)=0$ for all $i$, as such the graph $\Gamma$ has no edges and $A$ is the zero matrix. ∎
Proposition 2.8.
Suppose $A$ is an adjacency $n\times n$ matrix corresponding to a graph $\Gamma$ without multiple edges (but it possibly has loops), then $A^{2}=\mu A$ if and only if $\Gamma$ is a disjoint union of complete graphs with $\mu$ vertices, where $\mu$ is the scalar satisfying $A^{2}=\mu A$.
Proof.
Suppose there exists a scalar $\mu$ such that $A^{2}=\mu A$. As $\Gamma$ does not have multiple edges, $A$ is a symmetric matrix with entries being either 0 or $1$. Further, as we are allowing loops, we can have $A_{ii}\neq 0$. Similar to the proof of the previous proposition, we have
$\mu A_{ii}=(A^{2})_{ii}=\sum_{j=1}^{n}A_{ij}A_{ji}=\sum^{n}_{j=1}A_{ij}=\deg(i)$
where $\deg(i)$ is equal to the number of neighbors of $i$ (including loops). As $A^{2}=\mu A$, it follows $\deg(i)=0$, if $A_{ii}=0$, or $\deg(i)=\mu$, if $A_{ii}=1$. As such, we can partition our vertices of $A$ into a set of isolated vertices and the set $S=\{i:A_{ii}=1\}$, where every element $i$ of $S$ has $\deg(i)=\mu$.
Restricting our attention to off-diagonal entries, we have $(A^{2})_{ij}=\mu A_{ij}$, so if $A_{ij}=0$, then $A_{ij}^{2}=0$ and there are no $2$-paths from $i$ to $j$. If $A_{ij}=1$, then $(A^{2})_{ij}=\mu$, so there are $\mu$ common neighbors of $i$ and $j$. As a corollary, the induced subgraph on $S$ is a disjoint union of complete graphs (with loops), each with $\mu$ vertices. ∎
Theorem 2.9.
Suppose $A$ is an $n\times n$ adjacency matrix with $A^{2}=\mu A$.
  1. (1)
    If $\mu=0$, then $A^{2}=0$ and all vertices are isolated.
  2. (2)
    If $\mu\neq 0$, then $A$ is a block-diagonal matrix with blocks $A_{1},A_{2},\ldots A_{m}$ (and possibly zero rows/columns for isolated vertices). Each block $A_{i}$ can be described using an equation
    $A_{i}=a_{i}\cdot a_{i}^{T},\qquad a_{i}\in\mathbb{Z}^{n_{i}}.$
    With this in mind, each $A_{i}$ is a symmetric, nonzero matrix of rank one.
    Additionally, for each block $i$,
[…]
Conversely, any matrix constructed in such a way satisfies $A^{2}=\mu A$.
Proof.
Suppose $A^{2}=\mu A$. If $\mu\neq 0$ then we can define $P:=\frac{1}{\mu}A$, which is a symmetric matrix with the same rank as $A$. Additionally,
$P^{2}=(\frac{1}{\mu}A)^{2}=\frac{1}{\mu^{2}}A^{2}=\frac{\mu}{\mu^{2}}A=\frac{1}{\mu}A=P,$
so $P$ is a projection matrix and thus can only have eigenvalues of 0 or $1$.
Consider the set $S=\{(i,j):P_{ij}>0\}$. As $P^{2}=P$, we have
$P_{ij}=\sum_{k=1}^{n}P_{ik}P_{kj}$
so if $P_{ij}>0$, there exists some $k$ with $P_{ik}>0$ and $P_{kj}>0$, hence connectivity on the induced subgraph from $S$ induces an equivalence relation. We thus partition $S$ into distinct components $C_{1},C_{2},\ldots C_{m}$.
Restricting $P$ to the component $C_{l}$, the submatrix $P_{l}$ is symmetric, entry-wise positive and idempotent. By the Perron-Frobenius theorem, the spectral radius is $1$, so $P_{l}$ has exactly one eigenvalue that is $1$, while the other eigenvalues are zero. It follows that the rank of $P_{l}$ is one, as $P_{l}$ is symmetric, and thus diagonlizable. As $P_{l}$ is rank one and symmetric, we may write $P_{l}=w\cdot w^{T}$ for some vector $w$. Additionally, we can normalize $w$ so that w=1||w||=1.
Letting $a:=\sqrt{\mu}w$, it follows
$A_{l}=\mu P_{l}=\mu w\cdot w^{T}=a\cdot a^{T}.$
[…]
Conversely, any block $a\cdot a^{T}$ satisfies
$(a\cdot a^{T})^{2}=(a^{T}\cdot a)(a\cdot a^{T})=\mu(a\cdot a^{T})$
thus $A^{2}=\mu A$. ∎
Example 2.10.
Consider the integer vector $a=\begin{pmatrix}1\\ 2\end{pmatrix}$ and the corresponding matrix
$A=a\cdot a^{T}=\begin{pmatrix}1\\ 2\end{pmatrix}\begin{pmatrix}1&&2\end{pmatrix}=\begin{pmatrix}1&&2\\ 2&&4\end{pmatrix}$.
[…]
Corollary 2.11.
Let $Q$ be a finite quiver with symmetric adjacency matrix $A$ over a field $K$, and assume $A^{3}\neq 0$. Then $Q$ satisfies the quantum Yang–Baxter equation if and only if, after a suitable reordering of vertices, $A$ is block diagonal
$A\;=\;\operatorname{diag}(A_{1},\dots,A_{m},0,\dots,0),$
where each nonzero block $A_{i}$ corresponds to a connected component of $Q$ and can be written as
$A_{i}\;=\;a_{i}a_{i}^{\mathsf{T}},\qquad a_{i}\in K^{n_{i}},$
with all coordinates of $a_{i}$ nonzero.
In particular, each nontrivial connected component of $Q$ is a complete weighted quiver (that is, there is a nonzero weighted arrow between any two vertices in the component).
Proof.
By the computation in Section 2, a quiver $Q$ with adjacency matrix $A$ satisfies the quantum Yang–Baxter equation if and only if the matrix $X=A\otimes A$ satisfies
$(X\otimes I_{n})(I_{n}\otimes X)(X\otimes I_{n})\;=\;(I_{n}\otimes X)(X\otimes I_{n})(I_{n}\otimes X)$
in $\operatorname{End}(K^{n}\otimes K^{n}\otimes K^{n})$, and this is equivalent to the single matrix identity
(2) $A^{2}\otimes A^{3}\otimes A\;=\;A\otimes A^{3}\otimes A^{2}.$
By Theorem 2.3, and using the assumption $A^{3}\neq 0$, the equality (2) holds if and only if there exists a scalar $\mu\in K$ such that
$A^{2}\;=\;\mu A.$
Thus $Q$ satisfies QYBE if and only if its adjacency matrix $A$ satisfies the “Temperley–Lieb idempotent” relation $A^{2}=\mu A$. The structure of such matrices is described in Theorem 2.9: after reordering vertices, $A$ is block diagonal,
$A\;=\;\operatorname{diag}(A_{1},\dots,A_{m})$
(possibly with additional zero rows/columns for isolated vertices), and each nonzero block $A_{i}$ can be written as
$A_{i}\;=\;a_{i}a_{i}^{\mathsf{T}},\qquad a_{i}\in\mathbb{Z}^{n_{i}},$
so each $A_{i}$ is symmetric of rank one and satisfies $A_{i}^{2}=\mu_{i}A_{i}$ with $\mu_{i}=a_{i}^{\mathsf{T}}a_{i}$.
Now fix a nonzero block $A_{i}$ corresponding to a connected component $C_{i}$ of $Q$. We show that all coordinates of $a_{i}$ are nonzero. Write $a_{i}=(a_{i,1},\dots,a_{i,n_{i}})^{\mathsf{T}}$, and suppose, for a contradiction, that $a_{i,j}=0$ for some $j$. Partition the vertex set of $C_{i}$ as
$S:=\{k:a_{i,k}\neq 0\},\qquad T:=\{k:a_{i,k}=0\}.$
If $k\in S$ and $\ell\in T$, then
$(A_{i})_{k\ell}\;=\;a_{i,k}a_{i,\ell}\;=\;a_{i,k}\cdot 0\;=\;0,$
and similarly $(A_{i})_{\ell k}=0$. Thus there are no edges between $S$ and $T$. Moreover, if $k,\ell\in T$, then
$(A_{i})_{k\ell}\;=\;a_{i,k}a_{i,\ell}\;=\;0,$
so there are no edges within $T$ either. Hence every vertex in $T$ is isolated in the component $C_{i}$, and there are no edges between $S$ and $T$; this contradicts the assumption that $C_{i}$ is connected and $A_{i}\neq 0$. Therefore $T$ must be empty, and all coordinates of $a_{i}$ are nonzero.
It follows that for each nonzero block $A_{i}=a_{i}a_{i}^{\mathsf{T}}$ corresponding to a connected component, we have
$(A_{i})_{k\ell}\;=\;a_{i,k}a_{i,\ell}\;\neq\;0\quad\text{for all }1\leq k,\ell\leq n_{i},$
so there is a nonzero weighted edge between every pair of vertices in $C_{i}$. Thus each nontrivial connected component of $Q$ is a complete weighted quiver. Isolated vertices correspond to zero blocks and do not affect the QYBE condition.
This proves the “only if” direction.
Conversely, suppose that after reordering vertices, $A$ is block diagonal with nonzero blocks $A_{i}=a_{i}a_{i}^{\mathsf{T}}$ corresponding to connected components, where each $a_{i}$ has all entries nonzero, so that each component is a complete weighted quiver. Then for each $i$ we have
$A_{i}^{2}\;=\;(a_{i}a_{i}^{\mathsf{T}})(a_{i}a_{i}^{\mathsf{T}})\;=\;(a_{i}^{\mathsf{T}}a_{i})\,(a_{i}a_{i}^{\mathsf{T}})\;=\;\mu_{i}A_{i},$
where $\mu_{i}=a_{i}^{\mathsf{T}}a_{i}\in K$. Zero blocks (isolated vertices) satisfy $0^{2}=0$ trivially. Hence $A^{2}=\mu A$ holds blockwise (with appropriate scalars on each component), and in particular $A$ satisfies the matrix identity (2). By Theorem 2.3 this is equivalent to $X=A\otimes A$ satisfying the quantum Yang–Baxter equation. Therefore $Q$ satisfies QYBE.
This establishes the equivalence and completes the proof. ∎
Next, we proceed to show that a connected quiver with a groupoid structure satisfies the QYBE. Recall that a groupoid over a set $M$ is a set $G$ equipped with source and target mappings $\alpha,\beta:G\rightarrow M$, a multiplication map $m$ from $G_{2}=\{(g,h)\in G\times G:\beta(g)=\alpha(h)\}$ to $G$, an injective units mapping $\epsilon:M\rightarrow G$, and an inversion mapping $i:G\rightarrow G$, satisfying the following properties (where we write $gh$ for $m(g,h)$ and $g^{-1}$ for $i(g)$):
  1. (1)
    $g(hk)=(gh)k$.
  2. (2)
    $\epsilon(\alpha(g))g=g=g\epsilon(\beta(g))$.
  3. (3)
    $gg^{-1}=\epsilon(\alpha(g)0$ and $g^{-1}g=\epsilon(\beta(g))$.
Proposition 2.12.
Let $Q$ be a quiver arising from a groupoid with connected components $C_{1},C_{2},\ldots C_{k}$. Let $A$ be the adjacency matrix with $A_{C_{1}},A_{C_{2}},\ldots,A_{C_{k}}$ the block matrices corresponding to the connected components. With this set-up, $A_{C_{i}}\otimes A_{C_{i}}$ satisfies the QYBE for each $i$. Equivalently, there exists a scalar $\mu_{i}$ such that $A^{2}_{C_{i}}=\mu_{i}A_{C_{i}}$. Furthermore, $A^{2}=\mu A$ for a single scalar $\mu$ if and only if $\mu_{1}=\mu_{2}=\ldots=\mu_{k}$.
Proof.
Let $\Lambda$ be the object set of the groupoid $\mathcal{G}$ arising from $Q$. For our groupoid $\mathcal{G}$, let
$g_{x}:=|\text{Aut}(x)|\qquad(x\in\Lambda)$
If $i$ and $j$ lie in the same connected component, then $\text{Hom}(i,j)$ is nonempty and
$|\text{Hom}(i,j)|=|\text{Aut}(j)|$
thus
$A_{ij}=|\text{Hom}(i,j)|=g_{j}$
If $i$ and $j$ are in different components, then $A_{ij}=0$, so each connected component with objects indexed by $C_{i}:=\{i_{1},i_{2},\ldots,i_{n}\}$ produces a block
$A_{C_{i}}=\textbf{1}_{n}\cdot(g_{i_{1}},\ldots,g_{i_{n}}),$
where $\textbf{1}_{n}$ is the $n\times 1$ column vector consisting of ones. Thus, for any connected component $C$ the corresponding block $A_{C}$ is rank one with $A_{C}=\textbf{1}\cdot g^{T}$. Lastly,
$A_{C_{i}}^{2}=(\textbf{1}\cdot g^{T})(\textbf{1}\cdot g^{T})=\textbf{1}(g^{T}\cdot\textbf{1})g^{T}=(\sum_{x\in C_{i}}g_{x})\textbf{1}\cdot g^{T}=\mu_{i}A_{C_{i}}.$
As such, each connected component satisfies $A^{2}\otimes A^{3}\otimes A=A\otimes A^{3}\otimes A^{2}$ by Theorem 2.3. The observations regarding the global matrix $A$ follows immediately from the above observations.
Remark 2.13.
As illustrated in the proof of the previous proposition, the scalar $\mu_{i}$ with $A_{C_{i}}^{2}=\mu_{i}A_{C_{i}}$ is the sum of the automorphism-group sizes in the $C_{i}$ connected component.
Corollary 2.14.
Let $Q$ be a connected quiver with a groupoid structure, then $Q$ satisfies the quantum Yang-Baxter equation.
From the previous corollary, it follows every connected groupoid induces a bialgebra structure via the FRT construction.
Example 2.15.
Consider a quiver with two connected components with a single object in both and $g_{1}=1$ and $g_{2}=2$. The corresponding matrix is
$A=\begin{bmatrix}1&&0\\ 0&&2\end{bmatrix}$.
There is not a scalar $\mu$ such that $A^{2}=\mu A$, despite both connected components satisfying $A_{C_{i}}^{2}=\mu_{i}A$ for some scalar $\mu_{i}$.
3. Construction of Hecke $R$-matrices
It is known that one can produce constant solutions to the quantum Yang-Baxter equation through the Temperley-Lieb algebra $TL_{N}$ [3, 12]. The Temperley-Lieb algebra $(TL_{N})$ has N1N-1 generators $\{X_{1},\ldots,X_{N-1}\}$ with relations
$X_{k}^{2}=dX_{k}$
$X_{k}X_{j}X_{k}=X_{k}\qquad\qquad|j-k|=1$
$X_{j}X_{k}=X_{k}X_{j}\qquad\qquad|j-k|>1.$
Temperley-Lieb algebra played a central role in Vaughan Jones’ discovery of his new polynomial invariant of knots which triggered a plethora of mathematical developments relating knot theory, topological quantum field theory, and statistical physics.
Here we follow Kulish’s rank-one tensor construction [11] to create a generator $X$ of $TL_{N}$ that we can then use to form the Hecke $R$-matrix $\check{R}=qI+X$. The matrix $X$ can be viewed as a weighted adjacency matrix of the complete quiver with $n^{2}\times n^{2}$ vertices, which will have special combinatorial properties, as we will see later.
Before describing $X$ in terms of quivers, we first remind the reader of the aforementioned construction of $X$ introduced by Kulish [11]. The recipe is as follows:
Fix a local realization of the $TL_{N}$ algebra $\{1,X_{1},X_{2},\ldots,X_{N-1}\}$ in the space
$\mathcal{H}_{N}=\bigotimes_{k=1}^{N}\mathbb{C}^{n}_{k}.$
  1. (1)
    To define the above realization, take an $n\times n$ invertible matrix $b$, with inverse $\overline{b}=b^{-1}$.
  2. (2)
    Through vectorization, view $b,\overline{b}$ as vectors of the $n^{2}$ dimensional space $\mathbb{C}^{n}\otimes\mathbb{C}^{n}$ with components $b_{ij},i,j=1,2,\ldots,n$.
  3. (3)
    Define the rank-one, Temperley-Lieb idempotent matrix $X\in\text{End}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})$ as
    $\text{vec}(b)\otimes\text{vec}(\overline{b}),\qquad X_{ab;cd}=b_{ab}\overline{b}_{cd}.$
    $X$ is a generator of $TL_{N}$ with $X_{k}$ acting on $\mathbb{C}^{n}_{k}\otimes\mathbb{C}^{n}_{k+1}$ of $\mathcal{H}_{N}$.
  4. (4)
    The matrix $\check{R}=qI+X$ is a Hecke $R$-matrix satisfying the quantum Yang-Baxter equation
    $\check{R}_{12}\check{R}_{23}\check{R}_{12}=\check{R}_{23}\check{R}_{12}\check{R}_{23}$
    where $\check{R}_{12}=\check{R}\otimes I$ and $\check{R}_{23}=I\otimes\check{R}$.
Quiver Realizations of $X$
From the above construction, matrices of the form $X=\text{vec}(b)\otimes\text{vec}(\overline{b})$ lead to nontrivial $R$-Hecke matrices. Following in the vein of the adjacency matrices of the previous section, we ask which quivers have adjacency matrices that can be expressed as $\text{vec}(b)\otimes\text{vec}(\overline{b})$.
To begin, $b$ is an arbitrary $n\times n$ invertible matrix, as such, $\text{vec}(b)\,\otimes\,\text{vec}(\overline{b})$ can have negative entries. With this in mind, it is more natural to consider weighted quivers, otherwise, the restrictions we impose on $b$ will be too rigid to lead to interesting Hecke $R$-matrices. Regardless, the weighted quivers we obtain have an underlying quiver structure that is similar to the quivers we have looked at in previous section; the connected components of the quiver are complete.
The following theorem is a simple yet quite useful observation, but before stating the theorem, we introduce some terminology. We will call $b_{ij}$ the source entries and $\overline{b}_{kl}$ the target entries for reasons that will become apparent in the proof of the following theorem. Furthermore, as $b$ can be viewed as a weighted matrix, we may view any $b_{ij}$ as a weight $\alpha_{ij}$. We call such an $\alpha_{ij}$ a source weight. Similarly, we may view $\overline{b}_{kl}$ as a weight $\beta_{kl}$, which we call a target weight.
Theorem 3.1.
Let $X$ be a $n^{2}\times n^{2}$ matrix. Then $X=\text{vec}(b)\,\otimes\,\text{vec}(\overline{b})$ for $b$ an invertible $n\times n$ matrix if and only if $X$ is a weighted adjacency matrix of a quiver $Q_{X}$ with the following structure:
  • $Q_{X}$ has $n^{2}$ vertices indexed as pairs $\{(i,j):1\leq i,j\leq n\}$.
  • Edges exist between two vertices $(i,j)$ and $(k,l)$ exactly when the entries $b_{ij}$ and $\overline{b}_{kl}$ are both nonzero. In particular, $Q_{X}$ is a complete graph among the nonzero source entry vertices and the nonzero target entry vertices.
  • The weights of $Q_{X}$ factor multiplicatively as $w(\text{source})\cdot w(\text{target})$. Additionally, the source weights and the target weights form inverse matrices.
Lastly, we observe the following four-cycle equation holds:
$X_{(i,j),(k,l)}\cdot X_{(i^{\prime},j^{\prime}),(k^{\prime},l^{\prime})}=X_{(i,j),(k^{\prime},l^{\prime})}\cdot X_{(i^{\prime},j^{\prime}),(k,l)}$
Proof.
First, we suppose $X=vec(b)\,\otimes vec(\overline{b})$ for $b$ an invertible $n\times n$ matrix. We write $u=vec(b)$ and $v=vec(\overline{b})$, thus $X=uv^{T}$, where $u,v$ are column vectors with $n^{2}$ entries and $X$ is an $n^{2}\times n^{2}$ matrix.
Regarding the first bullet point, suppose we have a basis $\{e_{1},e_{2},\ldots,e_{n}\}$ for $\mathbb{C}^{n}$, then $\{e_{i}\otimes e_{j}\}_{i,j=1}^{n}$ is a basis of $\mathbb{C}^{n}\otimes\mathbb{C}^{n}$, meaning the basis vectors are indexed by ordered pairs $(i,j)$. In particular, any operator $X:\mathbb{C}^{n}\otimes\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}\otimes\mathbb{C}^{n}$ is represented by an $n^{2}\times n^{2}$ matrix whose rows and columns are indexed by pairs $(i,j)$. In the $\{e_{i}\otimes e_{j}\}_{i,j=1}^{n}$ basis, we may write $vec(b)=\displaystyle{\sum_{i,j}^{n}b_{i,j}e_{i}\otimes e_{j}}$ and $X_{(i,j),(k,l)}=b_{ij}\cdot\overline{b}_{kl}$. As the basis vectors of $\mathbb{C}^{n}\otimes\mathbb{C}^{n}$ are $e_{i}\otimes e_{j}$, the vertices of $Q_{X}$ can be expressed as
$\text{vertices}(Q_{X})=\{(i,j)\,\mid\,1\leq i,j\leq n\}.$
For the second bullet point, the entries of $X_{(i,j),(k,l)}$ are defined as
$X_{(i,j),(k,l)}=b_{ij}\overline{b}_{kl};$
therefore, if $b_{ij}=0$, then $X_{(i,j),(k,l)}=0$ for all $(k,l)$, meaning vertex $(i,j)$ has not outgoing arrows. Similarly, if $\overline{b}_{kl}=0$, then $(k,l)$ has no incoming arrows. This is the intuition behind calling $b_{ij}$ source entry vertices and $\overline{b}_{kl}$ target entry vertices. Thus in order to have an arrow $(i,j)\rightarrow(k,l)$, one must have both $b_{ij}\neq 0$ and $\overline{b}_{kl}\neq 0$.
The observation that $Q_{X}$ is complete among nonzero source entries $b_{ij}$ and nonzero target entries $\overline{b}_{kl}$ follows immediately from the above realization.
For the final bullet point, suppose we have an arrow $(i,j)\rightarrow(k,l)$, then if we label $(i,j)$ with the scalar $\alpha_{ij}=u(i,j)=b_{ij}$ and the vertex $(k,l)$ with the scalar $\beta_{k,l}=v(k,l)=\overline{b}_{kl}$, then the weight of the arrow $(i,j)\rightarrow(k,l)$ is
$w((i,j),(k,l))=X_{(i,j),(k,l)}=b_{ij}\cdot\overline{b}_{kl}=\alpha_{ij}\cdot\beta_{k,l}=w((i,j))\cdot w((k,l)).$
By definition, we have $b=(\alpha_{ij})$ and $\overline{b}=(\beta_{kl})$, meaning the source weights and the target weights form inverse matrices.
The converse of the theorem follows by defining $b=(\alpha_{ij})$ and $\overline{b}=(\beta_{kl})$; the structure is entirely determined by the weights.
The four-cycle identity is immediate from definition:
$X_{(i,j),(k,l)}\cdot X_{(i^{\prime},j^{\prime}),(k^{\prime},l^{\prime})}$ $=b_{ij}\cdot\overline{b}_{kl}\cdot b_{i^{\prime}j^{\prime}}\cdot\overline{b}_{k^{\prime}l^{\prime}}$
$=b_{ij}\cdot\overline{b}_{k^{\prime}l^{\prime}}\cdot b_{i^{\prime}j^{\prime}}\cdot\overline{b}_{kl}$
$=X_{(i,j),(k^{\prime},l^{\prime})}\cdot X_{(i^{\prime},j^{\prime}),(k,l)}.$
The four-cycle identity is really just equivalent to $X$ being a rank-one matrix. ∎
Example 3.2.
Consider the following weighted quiver:
The weighted adjacency matrix is given by
$X=\begin{pmatrix}-2&&1&&3/2&&-1/2\\ -4&&2&&3&&-1\\ -6&&3&&9/2&&-3/2\\ -8&&4&&6&&-2\end{pmatrix}$
which can be expressed as
$X=\text{vec}(b)\otimes\text{vec}(\overline{b})$
where
$b=\begin{pmatrix}1&&2\\ 3&&4\end{pmatrix},\qquad\qquad\overline{b}=-1/2\begin{pmatrix}4&&-2\\ -3&&1\end{pmatrix}$
As a result, the matrix $X$ is a Temperley-Lieb idempotent matrix and $\check{R}=qI+X$ is a Hecke $R$-matrix.
To demonstrate the four-cycle identity, we let $i=j=k=1,l=2$ and $i^{\prime}=k^{\prime}=l^{\prime}=1,j^{\prime}=2$. We then find
$X_{(1,1),(1,2)}\cdot X_{(1,2),(1,1)}$ $=X_{12}\cdot X_{21}$
$=(1)\cdot(-4)$
$=(-2)\cdot(2)$
$=X_{11}\cdot X_{22}$
$=X_{(1,1),(1,1)}\cdot X_{(1,2),(1,2)}$
The above equality is visualized below. The main point is that the product of the weights of the red arrows is equal to the product of the weights of the blue arrows.
Example 3.3.
Let $Q$ be the Toeplitz quiver below:
Let $V=\text{span}\{a,b\}$, so $V\otimes V=\text{span}\{a\otimes a,a\otimes b,b\otimes a,b\otimes b\}$.
Let $\psi=a\otimes b-b\otimes a\in V\otimes V$. A matrix representation for $\psi$ can be written as
$\psi]_{\mathcal{B}}=\begin{pmatrix}0\\ 1\\ -1\\ 0\end{pmatrix}$,
which is the vectorization of the invertible matrix
[…]
Therefore, we can compute the rank one matrix
$X=vec(b)\otimes vec(\overline{b})=\begin{pmatrix}0\\ 1\\ -1\\ 0\end{pmatrix}\otimes\begin{pmatrix}0\\ -1\\ 1\\ 0\end{pmatrix}=\begin{pmatrix}0\\ 1\\ -1\\ 0\end{pmatrix}\cdot\begin{pmatrix}0&-1&1&0\end{pmatrix}=\begin{pmatrix}0&&0&&0&&0\\ 0&&-1&&1&&0\\ 0&&1&&-1&&0\\ 0&&0&&0&&0\end{pmatrix}$
The matrix $X$ satisfies the Temperley-Lieb relations and thus can be used to define a Hecke $R$-matrix
$\check{R}(q)=qI+X,$
as well as other Hecke $R$-matrices (see for example [11]). The matrix $X$ could also be viewed as the weighted adjacency matrix of the complete weighted quiver
Construction involving a projective matrix $P$
Observe $X^{2}=2X$; therefore, in defining $P=X/2$, we have $P^{2}=P$. Now consider the matrix
$e=-(q+q^{-1})X=\alpha P,\qquad\text{where}\qquad\alpha=-2(q+q^{-1}).$
We want to determine whether $e$ satisfies the three Temperley-Lieb algebra relations:
  1. (1)
    $e_{i}^{2}=\alpha e_{i}$
  2. (2)
    $e_{i}e_{j}e_{i}=e_{i},\quad|i-j|=1$
  3. (3)
    $e_{i}e_{j}=e_{j}e_{i},\quad|i-j|>1$
where $e_{i}=I^{\otimes(i-1)}\times e\times I^{\otimes(n-i-1)},1\leq i\leq n-1$.
The first condition is satisfied:
$e^{2}=\alpha^{2}P^{2}=\alpha^{2}P=\alpha(\alpha P)=\alpha e,$
furthermore, the third condition is satisfied trivially, so this leaves us with proving the second condition.
To investigate the second condition, we first restrict our attention to $P$. A quick computation in Macaulay2 shows
$P_{12}P_{23}P_{12}=\frac{1}{4}P_{12}.$
thus
$e_{12}e_{23}e_{12}=\alpha^{3}P_{12}P_{23}P_{12}=\frac{\alpha^{3}}{4}P_{12}.$
For the second condition to be met, we must have $e_{12}=\frac{\alpha^{3}}{4}P_{12}$, which could be rewritten as $\alpha P_{12}=\frac{\alpha^{3}}{4}P_{12}$. Solving for $\alpha$ gives
$\alpha=\frac{\alpha^{3}}{4}\Rightarrow\alpha^{2}=4\Rightarrow\alpha=\pm 2.$
As $\alpha=-2(q+q^{-1})$, we must have
$\pm 2=-2(q+q^{-1})\Rightarrow\mp 1=q+q^{-1}\Rightarrow q^{2}\pm q+1=0$
which has solutions
$q=\frac{1\pm 3i}{2}=e^{\pm\pi i/3}\qquad\text{and}\qquad q=\frac{-1\pm 3i}{2}=e^{\pm 2\pi i/3}.$
In conclusion, the first and the third TL-algebra relations hold for all $q$; however, the second condition holds only when
$q\in\{e^{\pm\pi i/3},e^{\pm 2\pi i/3}\}.$
4. Application: Leavitt path algebra over quiver with relations
Let $Q=(Q_{0},Q_{1},s,r)$ be a quiver where $Q_{0}$ is the set of vertices, $Q_{1}$ is the set of arrows and both $s$, and $r$ are functions from $Q_{1}$ to $Q_{0}$ such that for any arrow $e$ from $u$ to $v$, $s(e)=u$ and $r(e)=v$ and let $\mathbb{K}$ be a field. A quiver with relations is a pair $(Q,R)$ where $Q$ is a quiver and $R$ is a two-sided ideal of $\mathbb{K}Q$ generated by a set of relations. The path algebra over a quiver with relations $(Q,R)$ is
$P_{\mathbb{K}}(Q,R)=\mathbb{K}Q/R.$
Given a quiver $Q$, we extend it to a new quiver $\overline{Q}$, where for each arrow $e\in Q_{1}$, we have, in addition, an arrow $e^{\ast}$ in the opposite direction, called the ghost arrow. So we have $r(e^{\ast})=s(e)$, and $s(e^{\ast})=r(e)$. We call the new quiver $\overline{Q}$ the double quiver. The Leavitt path algebra $L_{\mathbb{K}}(Q)$ over quiver $Q$ is defined as the quotient algebra of $\mathbb{K}\overline{Q}$ by the so-called Cuntz-Krieger relations.
(CK1) $e^{\ast}f=\delta_{e,f}r(e)$ for each $e,f\in Q_{1}$.
(CK2) $v=\sum_{e\in Q_{1},s(e)=v}ee^{\ast}$ for every regular vertex (that is, vertex which emits at least one and only finitely many arrows) $v\in Q_{0}$.
Now, suppose we have a quiver with relations $(Q,R)$, then the Leavitt path algebra over $(Q,R)$ is defined as
$L_{\mathbb{K}}(Q,R)=L_{\mathbb{K}}(Q)/R.$
This motivates us to study quivers with RTT relations determined by Hecke R-matrices and we study Leavitt path algebras over quivers with such relations.
Let $Q=(Q_{0},Q_{1},s,r)$ be a row-finite quiver. For each $u\in Q_{0}$, let $L(u):=\operatorname{span}_{\mathbb{K}}\{l\in Q_{1}:s(l)=r(l)=u\}$ be the $\mathbb{K}$-vector space of loops at $u$. Consider invertible linear maps $R_{u}\in GL(L(u)\otimes L(u))$ for each $u\in E_{0}$ satisfying the following conditions:
  • (1)
    (Quantum Yang-Baxter Equation)
    $(R_{u}\otimes I)(I\otimes R_{u})(R_{u}\otimes I)=(I\otimes R_{u})(R_{u}\otimes I)(I\otimes R_{u})$
  • (2)
    (Hecke Condition)
    $R_{u}^{2}=1+(q-q^{-1})R_{u}$
Let $L(u)$ be $k$-dimensional. Suppose $L(u)\cong U\otimes W$. We define the matrix of loop generators $X^{(u)}$ at $u$ as $X^{(u)}=\sum_{i,j}E_{ij}\otimes x_{ij}\in End(U)\otimes P_{\mathbb{K}}(Q,R)$. We impose the following RTT condition:
$R_{u}X_{1}^{(u)}X_{2}^{(u)}=X_{2}^{(u)}X_{1}^{(u)}R_{u}$
where $X_{1}^{(u)}=X^{(u)}\otimes I$ and $X_{2}^{(u)}=I\otimes X^{(u)}$.
Let $I_{loop}$ be the two-sided ideal of $\mathbb{K}Q$ generated by the union of all these local RTT relations over $u\in Q_{0}$.
The path algebra of the quiver with these RTT relations determined by Hecke condition is defined as $P_{\mathbb{K}}(Q,R)=\mathbb{K}Q/I_{loop}$ and the Leavitt path algebra over the quiver with these RTT relations determined by Hecke condition is defined as the quotient algebra of $P_{\mathbb{K}}(\overline{Q},R)$ modulo Cuntz-Krieger conditions.
We give below some examples to illustrate the above construction.
Example 4.1.
Consider a quiver $Q$ with single vertex $u$ and two loops $a$ and $b$.
$\bullet u$$\scriptstyle{a}$$\scriptstyle{b}$
Here $L(u)=span_{\mathbb{K}}\{a,b\}$. Fix the ordered basis $\{a\otimes a,a\otimes b,b\otimes a,b\otimes b\}$ of $L(u)\otimes L(u)$. Define the invertible linear operator $R_{u}$ on $L(u)\otimes L(u)$ as
$R_{u}=\left[{\begin{array}[]{cccc}q&0&0&0\\ 0&q-q^{-1}&1&0\\ 0&1&0&0\\ 0&0&0&q\end{array}}\right]$
We have $R_{u}(a\otimes a)=qa\otimes a$, $R_{u}(b\otimes b)=qb\otimes b$, $R_{u}(a\otimes b)=b\otimes a+(q-q^{-1})a\otimes b$, $R_{u}(b\otimes a)=a\otimes b$.
It is not difficult to verify that $R_{u}$ satisfies the quantum Yang-Baxter equation, and it also satisfies the Hecke condition.
Here the matrix of loop generators $X^{(u)}$ is a $2\times 2$ matrix
$X^{(u)}=\left[{\begin{array}[]{cc}a&0\\ 0&b\end{array}}\right]$
Note that $X_{1}^{(u)}=X^{(u)}\otimes 1$ and $X_{2}^{(u)}=1\otimes X^{(u)}$. Clearly,
$X_{1}^{(u)}=\left[{\begin{array}[]{cccc}a&0&0&0\\ 0&a&0&0\\ 0&0&b&0\\ 0&0&0&b\end{array}}\right]\qquad\text{ and }\qquad X_{2}^{(u)}=\left[{\begin{array}[]{cccc}a&0&0&0\\ 0&b&0&0\\ 0&0&a&0\\ 0&0&0&b\end{array}}\right]$
Thus, the RTT condition $R_{u}X_{1}^{(u)}X_{2}^{(u)}=X_{2}^{(u)}X_{1}^{(u)}R_{u}$ reduces to $ab=qba$.
Therefore, the path algebra over quiver with these relations $(Q,R)$ is given as
$P_{\mathbb{K}}(Q,R)=\frac{\mathbb{K}\langle a,b\rangle}{(ab-qba)}$
It follows the path algebra over the above quiver with RTT conditions determined by Hecke condition is the quantum polynomial algebra.
The Leavitt path algebra $L_{\mathbb{K}}(Q,R)$ is a $\mathbb{K}$-algebra generated by $\{u,a,b,a^{*},b^{*}\}$ subject to the following relations:
$v^{2}=v$, $vx=xv=x$ for any $x\in\{a,b,a^{*},b^{*}\}$
$va=a=av$, $vb=b=bv$, $va^{*}=a^{*}=a^{*}v$, $vb^{*}=b^{*}=b^{*}v$
$a^{*}a=v$, $b^{*}b=v$, $a^{*}b=0$, $b^{*}a=0$
$v=aa^{*}+bb^{*}$
$ab=qba$.
Example 4.2.
Let $E$ be the following quiver
[…]
where $E_{0}=\{u\}$, with $4$ loops at $u$ denoted by $x_{ij}$, $1\leq i\leq 2$, $1\leq j\leq 2$. Here $L(u)=\operatorname{span}_{\mathbb{K}}\{x_{ij}:1\leq i\leq 2,1\leq j\leq 2\}$. We have $X^{(u)}=[x_{ij}]$. Choose standard Hecke $R$-matrix as
$R_{u}=\left[{\begin{array}[]{cccc}q&0&0&0\\ 0&1&q-q^{-1}&0\\ 0&0&1&0\\ 0&0&0&q\end{array}}\right]$
Note that $X_{1}^{(u)}=X^{(u)}\otimes 1$ and $X_{2}^{(u)}=1\otimes X^{(u)}$. Clearly, $X_{1}^{(u)}=X^{(u)}\otimes 1=\left[{\begin{array}[]{cccc}x_{11}&0&x_{12}&0\\ 0&x_{11}&0&x_{12}\\ x_{21}&0&x_{22}&0\\ 0&x_{21}&0&x_{22}\end{array}}\right]$ and $X_{2}^{(u)}=1\otimes X^{(u)}=\left[{\begin{array}[]{cccc}x_{11}&x_{12}&0&0\\ x_{21}&x_{22}&0&0\\ 0&0&x_{11}&x_{12}\\ 0&0&x_{21}&x_{22}\end{array}}\right]$
Now the RTT relation $R_{u}X_{1}^{(u)}X_{2}^{(u)}=X_{2}^{(u)}X_{1}^{(u)}R_{u}$ gives us the following relations:
$x_{11}x_{12}=qx_{12}x_{11},x_{21}x_{22}=qx_{22}x_{21}$
$x_{11}x_{21}=qx_{21}x_{11},x_{12}x_{22}=qx_{22}x_{12}$
$x_{12}x_{21}=x_{21}x_{12}$
$x_{11}x_{22}-x_{22}x_{11}=(q-q^{-1})x_{12}x_{21}$
Let $A$ be the subalgebra of the Leavitt path algebra over quiver with relations $L_{\mathbb{K}}(Q,R)$, generated only by the arrows $x_{ij}$ (ignoring $v$ and ghost arrows). By the universal property of FRT algebras, we have that $A$ is isomorphic to the quantum Matrix algebra $\mathcal{O}_{q}(M_{2})$.
Now, we can generalize this idea for a general $n\times n$ quantum matrix algebra.
Example 4.3.
Fix an integer $n\geq 2$ and a field $K$ together with a parameter $q\in K^{\times}$. Let $Q$ be the quiver consisting of a single vertex $u$ and $n^{2}$ loops
$Q_{0}=\{u\},\qquad Q_{1}=\{x_{ij}:1\leq i,j\leq n\},$
where each $x_{ij}$ is a loop at $u$:
$s(x_{ij})=r(x_{ij})=u.$
The $K$-vector space of loops at $u$ is
$L(u)=\mathrm{span}_{K}\{x_{ij}:1\leq i,j\leq n\}.$
We identify $L(u)$ with $K^{n}\otimes K^{n}$ by sending $x_{ij}$ to $e_{i}\otimes f_{j}$, where $\{e_{1},\dots,e_{n}\}$ and $\{f_{1},\dots,f_{n}\}$ are standard bases of $K^{n}$. We collect the loop generators into an $n\times n$ matrix
$X(u)=T:=(t_{ij})_{1\leq i,j\leq n},\qquad t_{ij}:=x_{ij}.$
Let $\{E_{ij}\}_{1\leq i,j\leq n}$ denote the standard matrix units in $M_{n}(K)$. The standard $R$-matrix $R\in\operatorname{End}(K^{n}\otimes K^{n})$ for $GL_{q}(n)$ is defined by
(3) $R\;=\;q\sum_{i=1}^{n}E_{ii}\otimes E_{ii}\;+\;\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}E_{ii}\otimes E_{jj}\;+\;(q-q^{-1})\sum_{\begin{subarray}{c}i,j=1\\ i>j\end{subarray}}^{n}E_{ij}\otimes E_{ji}.$
Equivalently, in terms of its action on the standard basis $\{e_{i}\otimes e_{j}\}_{1\leq i,j\leq n}$ of $K^{n}\otimes K^{n}$, one has
$R(e_{i}\otimes e_{j})=\begin{cases}q\,e_{i}\otimes e_{i},&i=j,\\[4.0pt] e_{i}\otimes e_{j},&i<j,\\[4.0pt] e_{i}\otimes e_{j}+(q-q^{-1})e_{j}\otimes e_{i},&i>j.\end{cases}$
It is well known that $R$ satisfies the quantum Yang–Baxter equation and the Hecke condition
We now define
$X_{1}:=T\otimes I_{n}\in\operatorname{End}(K^{n}\otimes K^{n})\otimes P_{\mathbb{K}}(Q,R),\qquad X_{2}:=I_{n}\otimes T\in\operatorname{End}(K^{n}\otimes K^{n})\otimes P_{\mathbb{K}}(Q,R),$
where $P_{\mathbb{K}}(Q,R)$ denotes the path algebra of the quiver with relations (to be imposed below), and $I_{n}$ is the $n\times n$ identity matrix.
At the vertex $u$ we impose the RTT relation
(4) $RX_{1}X_{2}=X_{2}X_{1}R.$
By definition, $P_{\mathbb{K}}(Q,R)$ is the path algebra of $Q$ modulo the two-sided ideal generated by the entries of the matrix equation (4).
Let us write $R$ in components as
$R(e_{i}\otimes e_{j})=\sum_{a,b=1}^{n}R_{ij}^{ab}\,e_{a}\otimes e_{b},$
so that $R_{ij}^{ab}\in K$ are the structure constants of $R$. Then the matrix equation (4) is equivalent to
(5) $\sum_{a,b=1}^{n}R_{ij}^{ab}\,t_{ak}t_{b\ell}\;=\;\sum_{a,b=1}^{n}t_{jb}t_{ia}\,R_{ab}^{k\ell},\qquad 1\leq i,j,k,\ell\leq n.$
Using the explicit form (3) of $R$, one can compute the nonzero components $R_{ij}^{ab}$:
$R_{ij}^{ab}=\begin{cases}q,&a=i=j,\,b=i,\\[4.0pt] 1,&a=i,\ b=j,\ i\neq j,\\[4.0pt] (q-q^{-1}),&a=i,\ b=j,\ i>j,\\[4.0pt] 0,&\text{otherwise},\end{cases}$
with the understanding that the $(q-q^{-1})$ term arises from the summands $E_{ij}\otimes E_{ji}$ for $i>j$.
Substituting these coefficients into (5) and organizing the resulting equations by cases on $(i,k)$, $(j,\ell)$, one arrives at the standard defining relations for the quantum matrix algebra: for all $1\leq i<k\leq n$ and $1\leq j<\ell\leq n$, the generators $t_{ij}$ satisfy
(6) $t_{ij}t_{i\ell}$ $=q\,t_{i\ell}t_{ij},$ (same row, increasing column)
(7) $t_{ij}t_{kj}$ $=q\,t_{kj}t_{ij},$ (same column, increasing row)
(8) $t_{ij}t_{k\ell}$ $=t_{k\ell}t_{ij},$ $\text{if }i<k\text{ and }j>\ell,$
(9) $t_{ij}t_{k\ell}-t_{k\ell}t_{ij}$ $=(q-q^{-1})t_{i\ell}t_{kj},$ $\text{if }i<k\text{ and }j<\ell.$
Relations (6)–(9) are precisely the Faddeev–Reshetikhin–Takhtajan (FRT) relations for the coordinate algebra of quantum matrices.
Now let $A$ be the subalgebra of the Leavitt path algebra $L_{\mathbb{K}}(Q,R)$ generated by the arrows $\{x_{ij}\}$, i.e.
[…]
ignoring the vertex idempotent and ghost arrows. By the FRT construction for $GL_{q}(n)$, the algebra presented by generators $\{t_{ij}\}$ and relations (6)–(9) is isomorphic to the standard quantum matrix algebra:
$A\;\cong\;\mathcal{O}_{q}(M_{n}).$
Thus, for the quiver $Q$ with one vertex and $n^{2}$ loops equipped with the RTT relations (4) determined by the Hecke $R$-matrix (3), the corresponding Leavitt path algebra $L_{\mathbb{K}}(Q,R)$ contains the quantum matrix algebra $\mathcal{O}_{q}(M_{n})$ as a canonical subalgebra generated by the loops $\{x_{ij}\}_{1\leq i,j\leq n}$.
Theorem 4.4.
Let $Q$ be a finite quiver with symmetric adjacency matrix $A$ satisfying $A^{2}=\mu A$ and $A^{3}\neq 0$. Then $P=\mu^{-1}A$ is an idempotent commuting with the flip $F$ on any tensor realization of $A$ (such as $A\otimes I$, $I\otimes A$, and linear combinations thereof), and $R(q)=qP-q^{-1}(I-P)$ solves the quantum Yang-Baxter equation and the Hecke condition. For each vertex $u$, the restriction of $R(q)$ to $L(u)\otimes L(u)$ yields Hecke matrices $R_{u}$ that produce the RTT relations of this section.
Proof.
Let $Q$ be a finite quiver with symmetric adjacency matrix $A$ satisfying $A^{2}=\mu A$ and $A^{3}\neq 0$ for some scalar $\mu\in K^{\times}$.
Define
$P:=\mu^{-1}A.$
Then
$P^{2}\;=\;\mu^{-2}A^{2}\;=\;\mu^{-2}\mu A\;=\;\mu^{-1}A\;=\;P,$
so $P$ is an idempotent. Since $A$ is symmetric, $A^{T}=A$, hence $P^{T}=P$, so $P$ is also symmetric.
Let $V$ be the vertex space with basis indexed by $Q_{0}$, and consider the tensor product $V\otimes V$. Let $F:V\otimes V\to V\otimes V$ be the flip map, $F(v\otimes w)=w\otimes v$. Any tensor realization of $A$ (such as $A\otimes I$, $I\otimes A$, and linear combinations thereof) is represented by a symmetric matrix; thus conjugating by $F$ simply transposes the corresponding matrix. As $P$ is symmetric, we have
$FPF=P^{T}=P.$
Equivalently, $P$ commutes with $F$:
$FP=PF$
on any tensor realization of $A$.
Define
$R(q):=qP\;-\;q^{-1}(I-P)\;=\;(q+q^{-1})P\;-\;q^{-1}I$
as an operator on the relevant tensor space ($V\otimes V$ or $L(u)\otimes L(u)$). We claim that $R(q)$ satisfies the Hecke relation
$(R(q)-qI)(R(q)+q^{-1}I)=0.$
A direct computation using only $P^{2}=P$ shows this. Indeed,
$R(q)-qI=qP-q^{-1}(I-P)-qI=qP-q^{-1}I+q^{-1}P-qI=(q+q^{-1})P-(q+q^{-1})I=(q+q^{-1})(P-I),$
and
$R(q)+q^{-1}I=qP-q^{-1}(I-P)+q^{-1}I=qP-q^{-1}I+q^{-1}P+q^{-1}I=(q+q^{-1})P.$
Thus
$(R(q)-qI)(R(q)+q^{-1}I)=(q+q^{-1})^{2}(P-I)P=(q+q^{-1})^{2}(P^{2}-P)=(q+q^{-1})^{2}(P-P)=0,$
since $P^{2}=P$. This proves the Hecke condition
$(R(q)-qI)(R(q)+q^{-1}I)=0.$
We now show that $R(q)$ satisfies the quantum Yang–Baxter equation
$R_{12}(q)R_{23}(q)R_{12}(q)=R_{23}(q)R_{12}(q)R_{23}(q),$
where $R_{12}(q)=R(q)\otimes I$ and $R_{23}(q)=I\otimes R(q)$ act on $V^{\otimes 3}$ (or, locally, on $L(u)^{\otimes 3}$).
First, note that $A^{2}=\mu A$ and $A$ symmetric imply, by Theorem 2.9, that after reordering vertices, $A$ is block diagonal with symmetric rank-one blocks:
$A=\operatorname{diag}(A_{1},\dots,A_{m}),\qquad A_{i}=a_{i}a_{i}^{\mathsf{T}},\quad a_{i}\in K^{n_{i}}.$
Consequently,
$P=\mu^{-1}A=\operatorname{diag}(P_{1},\dots,P_{m}),\qquad P_{i}=\mu_{i}^{-1}A_{i}$
is a direct sum of rank-one idempotents. Each $P_{i}$ is a projection onto the one-dimensional subspace $Ka_{i}$, and the corresponding tensor operators
$P_{i,12}:=P_{i}\otimes I,\qquad P_{i,23}:=I\otimes P_{i}$
on $K^{n_{i}}\otimes K^{n_{i}}\otimes K^{n_{i}}$ generate a representation of the Temperley–Lieb algebra:
$P_{i,12}^{2}=P_{i,12},\quad P_{i,23}^{2}=P_{i,23},\quad P_{i,12}P_{i,23}P_{i,12}=\frac{1}{\mu}P_{i,12},\quad P_{i,23}P_{i,12}P_{i,23}=\frac{1}{\mu}P_{i,23}.$
The last two relations follow from the rank-one nature of $P_{i}$: $P_{i,12}$ and $P_{i,23}$ are projections whose composition factors through the one-dimensional subspace $Ka_{i}\otimes a_{i}\otimes a_{i}$, and a straightforward computation shows that $P_{i,12}P_{i,23}P_{i,12}$ and $P_{i,23}P_{i,12}P_{i,23}$ reduce to $P_{i,12}$ and $P_{i,23}$, respectively.
By summing over the blocks, we obtain operators
$P_{12}:=P\otimes I,\qquad P_{23}:=I\otimes P$
on $V^{\otimes 3}$ satisfying the Temperley–Lieb relations
$P_{12}^{2}=P_{12},\quad P_{23}^{2}=P_{23},\quad P_{12}P_{23}P_{12}=\frac{1}{\mu}P_{12},\quad P_{23}P_{12}P_{23}=\frac{1}{\mu}P_{23}.$
Now set
$R_{12}(q):=qP_{12}-q^{-1}(I-P_{12}),\qquad R_{23}(q):=qP_{23}-q^{-1}(I-P_{23}).$
Using only the idempotent relations $P_{12}^{2}=P_{12}$, $P_{23}^{2}=P_{23}$ and the Temperley–Lieb relations $P_{12}P_{23}P_{12}=\frac{1}{\mu}P_{12}$, $P_{23}P_{12}P_{23}=\frac{1}{\mu}P_{23}$, a routine algebraic expansion (see, for example, Kulish [11]) shows that
$R_{12}(q)R_{23}(q)R_{12}(q)=R_{23}(q)R_{12}(q)R_{23}(q)$
as operators on $V^{\otimes 3}$. In other words, $R(q)$ solves the quantum Yang–Baxter equation.
For each vertex $u\in Q_{0}$, let $L(u)$ denote the $K$-vector space spanned by loops at $u$. The adjacency matrix $A$ decomposes compatibly with the direct sum $\bigoplus_{u\in Q_{0}}L(u)$, and the above construction restricts to each $L(u)\otimes L(u)$. Thus we obtain local operators
$R_{u}(q):=R(q)\big|_{L(u)\otimes L(u)}$
satisfying the same Hecke and the quantum Yang–Baxter relations. These $R_{u}(q)$ are the Hecke matrices that produce the RTT relations used to define the corresponding path and Leavitt path algebras over quivers with RTT relations. This completes the proof. ∎
Example 4.5.
Consider the quiver $Q$ below:
The quiver $Q$ has a groupoid structure with adjacency matrix given by
$A=\begin{pmatrix}2&&0&&0&&0\\ 0&&1&&1&&0\\ 0&&1&&1&&0\\ 0&&0&&0&&2\end{pmatrix}$
which satisfies the equation $A^{2}=2A$. As such, we define $P=\frac{1}{2}A$, which is an idempotent matrix. The flip matrix $F$ is given by
$F=\begin{pmatrix}1&&0&&0&&0\\ 0&&0&&1&&0\\ 0&&1&&0&&0\\ 0&&0&&0&&1\end{pmatrix}$.
It’s worth noting we have the relation $P=\frac{1}{2}(I+F)$. One can quickly check that FPPF=0FP-PF=0 (or just note $A$ is symmetric). With the three conditions outlined above satisfied, we may now define
$R(q)$ $=qP-q^{-1}(I-P)$
$=q(\frac{1}{2}(I+F))-q^{-1}(I-\frac{1}{2}(I+F))$
$=\frac{q}{2}I+\frac{q}{2}F-q^{-1}I+\frac{q^{-1}}{2}I+\frac{q^{-1}}{2}F$
$=\frac{q-q^{-1}}{2}I+\frac{q+q^{-1}}{2}F$
Letting ${a=\frac{q-q^{-1}}{2}}$ and ${b=\frac{q+q^{-1}}{2}}$, we may write $R(q)=aI+bF$, which is the matrix
$R(q)=\begin{pmatrix}a+b&&0&&0&&0\\ 0&&a&&b&&0\\ 0&&b&&a&&0\\ 0&&0&&0&&a+b\end{pmatrix}$.
This is a “classical” example and the matrix is known to satisfy the QYBE and the Hecke condition.